SUBMILLIMETER-WAVE ROTATIONAL SPECTRA OF DNC

T. AMANO, Department of Chemistry and Department of Physics and Astronomy, University of Waterloo, Waterloo, ON N2L 3G1, Canada.

Spectroscopic investigations of DNC have been less extensive than those for HNC. See Brünken et al.^{a} and Bechtel et al.^{b} for relevant references. In the present investigation, rotational transitions of DNC have been observed in the submillimeterwave region in an extended negative glow discharge in a gas mixture of CD_{4} and N_{2}. The dissociative recombination reaction of DCND^{+} with electrons is thought to be a dominant channel to produce DNC in highly excited vibrational states; the rotational lines in levels up to (008) are observed. The rotational and centrifugal distortion constants are determined for these states along with those for the (100) state. The measurement accuracy is high enough to determine some higher order vibration-rotation interaction constants. The least-squares fits were straightforward except for (004), (006), and (008), where very small but significant perturbations were found. A striking isotope effect was observed on the vibrational temperature in this investigation. The vibrational temperature for the ν_{3} mode for DNC is as high as 4000 K and the rotational transitions are observable in states up to (008), while the corresponding temperature is about 1500 K for HNC. The vibrational temperature for the ν_{1} mode is about 1000 K for DNC and about 1300 K for HNC. The bending vibrational mode is not excited, and the vibrational temperature for the ν_{2} mode is only about 400 K. The origin of this conspicuous excitation of the ν_{3} mode of DNC is not obvious. However, it should be closely related to mechanism of the dissociation of HCNH and DCND. Apparently the difference in the masses of the departing H/D is a factor causing this difference, but the vibrational temperature for ν_{3} of DCN is not particularly high, about 1000 K. When the D atom departs from the D-C side, apparently the C-N vibration is highly excited. On the other hand, when the D-N bond is broken, not much excitation of the C-N vibration occurs.

aS. Brünken, H. S. P. Müller, S. Thorwirth, F. Lewen, and G. Winnewisser, J. Mol. Strunct., 780-781,3 (2006)

bH. A. Bechtel, A. H. Steeves, and R. W. Field, Astrophys. J., 649,L53 (2006)